Silver chromate is sparingly soluble in aqueous solutions. The Kşp of Ag2CrO4 is 1.12 x 10-12 M3

What is the solubility (in moles per liter) of silver chromate in a 1.00 M potassium chromate aqueous solution? 

What is the solubility (in moles per liter) of silver chromate in a 1.00 M silver nitrate aqueous solution?

 What is the solubility (in moles per liter) of silver chromate in pure water? 
Answer #1

1)

K2CrO4 here is Strong electrolyte

It will dissociate completely to give [CrO42-] = 1.0 M

At equilibrium:

Ag2CrO4 <----> 2 Ag+ + CrO42-

   2s 1 + s

Ksp = [Ag+]^2[CrO42-]

1.12*10^-12=(2s)^2*(1+ s)

Since Ksp is small, s can be ignored as compared to 1

Above expression thus becomes:

1.12*10^-12=(2s)^2*(1)

1.12*10^-12=4*S^2

s = 5.292*10^-7 M

Answer: 5.29*10^-7 M

2)

AgNO3 here is Strong electrolyte

It will dissociate completely to give [Ag+] = 1.0 M

At equilibrium:

Ag2CrO4 <----> 2 Ag+ + CrO42-

   1 +2s s

Ksp = [Ag+]^2[CrO42-]

1.12*10^-12=(1 + 2 s)^2*(s)

Since Ksp is small, s can be ignored as compared to 1

Above expression thus becomes:

1.12*10^-12=(1)^2*(s)

1.12*10^-12= 1 * 1(s)^1

s = 1.12*10^-12 M

Answer: 1.12*10^-12 M

3)

At equilibrium:

Ag2CrO4 <----> 2 Ag+ + CrO42-

   2s s

Ksp = [Ag+]^2[CrO42-]

1.12*10^-12=(2s)^2*(s)

1.12*10^-12= 4(s)^3

s = 6.542*10^-5 M

Answer: 6.54*10^-5 M